# Sensor-fusion based Prognostics Framework for Complex Engineering Systems Exhibiting Multiple Failure Modes

### v1

@@ Abstract
Complex engineering systems are often subject to multiple failure modes. Developing a remaining useful life (RUL) prediction model that does not consider the failure mode causing degradation is likely to result in inaccurate predictions. However, distinguishing between causes of failure without manually inspecting the system is nontrivial. This challenge is increased when the causes of historically observed failures are unknown. Sensors, which are useful for monitoring the state-of-health of systems, can also be used for distinguishing between multiple failure modes as the presence of multiple failure modes results in discriminatory behavior of the sensor signals. When systems are equipped with multiple sensors, some sensors may exhibit behavior correlated with degradation, while other sensors do not. Furthermore, which sensors exhibit this behavior may differ for each failure mode. In this paper, we present a simultaneous clustering and sensor selection approach for unlabeled training datasets of systems exhibiting multiple failure modes. The cluster assignments and the selected sensors are then utilized in real-time to first diagnose the active failure mode and then to predict the system RUL. We validate the complete pipeline of the methodology using a simulated dataset of systems exhibiting two failure modes and on a turbofan degradation dataset from NASA.
Index Terms:
Multiple failure modes, Sensor-fusion, prognostics, Sensor Selection, Adaptive Sparse Group Lasso, Mixture of Gaussian regressions, Functional Principal Component Analysis
publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
I Introduction
@@
  • Finally, we provide a two-step sequential approach to leverage this knowledge of failure labels and informative sensors to (1) perform active failure mode diagnosis and (2) subsequent RUL prediction on systems degrading in real-time. The paper mentions this as the “online diagnostics and prognostics” step.
The remainder of the paper is outlined as follows. In Section II, we review the relevant literature associated with this research area. In Section III, we outline our proposed multi-failure mode prognostics framework. In Section IV, we present our offline sensor selection scheme that utilizes an EM algorithm. In Section V, we present our online diagnostics and prognostics method. It discusses a strategy to fuse informative sensors for real-time diagnosis of failure modes and subsequent prediction of RUL. Our methodology is demonstrated through two case studies – one using a simulated dataset in section VI and the other NASA turbofan engine dataset in section VII. We conclude by summarizing our framework and key findings in Section VIII.
II Literature Review
@@
III Overview of the Prognostics Framework
@@
IV Offline Sensor Selection Algorithm
@@
IV-A Feature Extraction: CA-FPCA
@@
IV-B Optimization of the MGR-ASGL model
@@
IV-B1 Mixture of Gaussian Regression (MGR) model
@@
$\begin{split}f_{\vartheta}(y_{i}\mid\boldsymbol{x}_{i,1},&\ldots,\boldsymbol{x}_{i,P})=\sum_{k=1}^{K}\pi_{k}\frac{1}{\sqrt{2\pi}\sigma_{k}}\times\\ &\ \exp\left(-\frac{1}{2\sigma_{k}^{2}}\left(y_{i}-\beta_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\beta}_{p,k}\right)^{2}\right)\end{split}$ (3)
IV-B2 Log-Likelihood
@@
$\begin{split}\ell(\boldsymbol{\Theta};&\boldsymbol{Y})=\sum_{i=1}^{N}\ln\sum_{k=1}^{K}\pi_{k}\times\\ &\ \frac{\rho_{k}}{\sqrt{2\pi}}\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)\end{split}\vspace*{-1em}$ (4)
@@
$\begin{split}\ell_{C}(\boldsymbol{\Theta};&\boldsymbol{Y},\mathbb{Z})=\sum_{i=1}^{N}\sum_{k=1}^{K}\boldsymbol{Z}_{i}[k]\ln\pi_{k}\left(\frac{\rho_{k}}{\sqrt{2\pi}}\right)\\ &\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)\end{split}\vspace*{-1em}$ (5)
@@
/* Feature Extraction using CA-FPCA from all $P$ sensors */
for $p=1$ to $P$ do
       $\mathcal{I}_{k}^{p}\leftarrow$ K-means clustering on FPC-scores;
       for $k=1$ to $K$ do
             $\xi_{i,m,p}(k)\leftarrow$ Use $\mathcal{I}_{k}^{p}$ to obtain CA-FPC-scores (Eq. 2);
       end for
      $x_{i,p}\leftarrow$ Obtain standardized CA-FPC-scores;
       $q_{p}\leftarrow$ Number of CA-FPC-scores retained for sensor $p$;
end for
/* MGR-ASGL based Expectation-Maximization (EM) algorithm */
Initialize $\mathcal{P}_{k}\leftarrow\emptyset$ for all $k$;
Initialize $\Psi^{0}\leftarrow 0,\Pi^{0}\sim\texttt{Uniform}(K),\boldsymbol{\rho}^{0}\leftarrow\texttt{randint}$;
$iter\leftarrow 0$;
while $iter\leq iter_{max}$ do
       $\gamma_{i}^{iter}\leftarrow$ Use $x_{i,p}$ in Eq. (7) for all $i$ ;
@@
       $\Pi^{iter}\leftarrow$ Update using Eq. (9) ;
@@
       $\boldsymbol{\rho}^{iter},\Psi^{iter}\leftarrow$ Solve Eq. (8) using $\sum_{p=1}^{P}\sqrt{q_{p}}$ and $\gamma_{i}^{iter}$;
       $iter\leftarrow iter+1$;
end while
$\gamma_{i}^{iter_{max}}\leftarrow$ Final responsibilities from Eq. (7);
$\boldsymbol{Z}_{i}[\operatorname*{arg\,max}_{k}{\gamma_{i}^{iter_{max}}}]\leftarrow 1$ for all $i$;
$\boldsymbol{Z}_{i}[\sim\operatorname*{arg\,max}_{k}{\gamma_{i}^{iter_{max}}}]\leftarrow 0$ for all $i$;
foreach $p$ and $k$ do
       if $\lVert\psi^{iter_{max}}_{p,k}\rVert_{2}\geq\delta$ then
@@
IV-B3 Adaptive Sparse Group Lasso (ASGL) penalty
@@
$\begin{split}\min_{\boldsymbol{\Theta}}-\frac{1}{N}\mathbb{E}_{\boldsymbol{\Theta}_{\text{old}}}[\ell_{C}(\boldsymbol{\Theta};&\boldsymbol{Y},\mathbb{Z})|\boldsymbol{Y}]+\lambda\sum_{k=1}^{K}\pi_{k}\sum_{p=1}^{P}\sqrt{q_{p}}\times\\ &\left[(1-\alpha)\left\|\boldsymbol{\psi}_{p,k}\right\|_{2}+\alpha\left\|\boldsymbol{\psi}_{p,k}\right\|_{1}\right]\end{split}$ (6)
@@
IV-B4 Expectation Maximization algorithm
@@
$\begin{split}\min_{\boldsymbol{\Theta}}\frac{1}{N}Q(\boldsymbol{\Theta}\mid&\boldsymbol{\Theta}_{\text{old}})+\lambda\sum_{k=1}^{K}\pi_{k}\sum_{p=1}^{P}\sqrt{q_{p}}\times\\ &\left[(1-\alpha)\left\|\boldsymbol{\psi}_{p,k}\right\|_{2}+\alpha\left\|\boldsymbol{\psi}_{p,k}\right\|_{1}\right]\end{split}\vspace*{-1em}$ (8)
@@
V Online diagnosis and Prognostics
@@
$\zeta_{(i),h}\leftarrow$ MFPCA on training units (Eq. 10);
Train a KNN classifier using $\zeta_{(i),h}$;
@@
$\zeta_{test,h}\leftarrow$ MFPCA on test unit;
$k^{*}\leftarrow$ Mode of labels in KNN of $\zeta_{test,h}$;
@@
$\zeta^{k^{*}}_{(i),h}\leftarrow$ MFPCA: selected sensors (Eq. 14);
$c_{0}^{k^{*}},\dots,c_{h}^{k^{*}}\leftarrow$ Obtain $\operatorname*{arg\,min}$ in Eq. 17;
@@
$\zeta^{k^{*}}_{test,h}\leftarrow$ MFPCA on test unit (Eq. 14);
$RUL_{test}\leftarrow$ Equation 18;
@@
V-A Sensor Fusion: MFPCA
@@
V-B Failure Mode Diagnosis: K-nearest Neighbors
@@
V-C RUL Prediction: Weighted Functional Regression
@@
V-C1 Time-Varying Functional Regression
@@
$\begin{split}\min_{c_{0}^{k^{*}},c_{1}^{k^{*}},\ldots,c_{H_{k^{*}}}^{k^{*}}}\sum_{i\in\mathcal{I}_{t^{*}}^{k^{*}}}&\left(\sqrt{w_{i}}y_{i}-\sqrt{w_{i}}c_{0}^{k^{*}}-\sum_{h=1}^{H_{k^{*}}}\sqrt{w_{i}}\zeta_{i,h}^{k^{*}}c_{h}^{k^{*}}\right)^{2}\\ &+\lambda\sum_{h=1}^{H_{k^{*}}}\lVert c_{h}^{k^{*}}\rVert\end{split}\vspace*{-0.5em}$ (17)
V-C2 RUL prediction
@@
VI Case Study 1: Simulated Dataset
@@ VI-A Simulation Model
Figure 2: Case study 1 data: Signals in sensors 1, 3, 5, 19 for both failure modes 1 and 2 under three different SNRs i.e., [2, 5], [5, 8], and [8, 11]
VI-B Failure Mode Labeling and Sensor Selection
We analyze the ability of the MGR-ASGL model to accurately cluster the degradation signals into one of the two possible failure modes along with selecting the subset of informative sensors. We start by utilizing CA-FPCA as mentioned in IV-A to extract features for each sensor separately. To determine the number of CA-FPC scores to retain, we use the 95% fraction of variance explained (FVE) criterion. For each cluster, we select the first $q_{p}^{k}$ CA-FPC-scores whose eigenvalues sum to at least 0.95. Since the number of CA FPC scores to retain can vary for each cluster, we retain the maximum of $q_{p}^{k}$ over both failure modes. For the remainder of this paper, we use 95% FVE criterion for determining the number of FPC scores, CA-FPC scores, and MFPC scores to retain. The CA-FPC scores serve as inputs to the MGR-ASGL model. Since this model has two tuning parameters, we utilize 5-fold cross-validation with $\lambda\in\{0.0050,0.0258,0.0466,...,0.3792,0.4000\}$ and $\alpha\in\{0.00,0.25,0.50,0.75,1.00\}$. We use minimum mean-squared error (minMSE) criterion, which selects the tuning parameters that minimize the mean-squared prediction error averaged over all folds. Table I displays the tuning parameter selection using both criteria over all three datasets. For each dataset and criterion, we fit the MGR-ASGL model using these tuning parameters.
TABLE I: Tuning parameter selection for case study 1
@@ VI-B Failure Mode Labeling and Sensor Selection
We then evaluate the performance of the MGR-ASGL model on the three datasets. While clustering is an unsupervised learning task, we know the ground truth regarding the failure modes. Therefore, we evaluate the ability of the clustering algorithm to match the ground truth. Furthermore, we have predetermined which sensors are most informative and we would like to determine whether or not the MGR-ASGL selects these sensors. The results of this performance evaluation are shown in Table II.
TABLE II: MGR-ASGL performance for case study 1
@@
For all scenarios, the initial failure mode labels were determined randomly. Table II shows that for the lowest and the highest signal-to-noise ratios, the clustering algorithm was capable of achieving at worst 78.8% accuracy for a particular failure mode. The exception is the dataset [5,8], where the algorithm has low accuracy at clustering. This is due to the tuning parameter combination resulting in minimal coefficient shrinkage that enabled the nonsignificant sensors to influence cluster assignments. A slightly larger $\lambda$ value would decrease this error. For all three models, several sensors that are deemed ”noninformative” were selected. This can be attributed to some of these sensors having high values of $\rho_{p}^{k}$ relative to the specified range of [0.1,0.6]. However, simply reporting the sensors selected does not account for how informative they are. Table III reports the four sensors with the highest $l_{2}$-norm of the regression coefficients associated with that sensor. Except for the [5,8] dataset, at least three of the four sensors with the largest $l_{2}$-norm for each failure mode are elements of the set of informative sensors. This indicates that the MGR-ASGL model results in the informative sensors having larger regression coefficients (in magnitude) than the noninformative sensors. Given the selected sensors, we want to test the robustness of the online portion of the methodology to predict remaining life.
TABLE III: $l_{2}$-norms of four most informative sensors for FM1 & FM2

### v2

@@ Abstract
Complex engineering systems are often subject to multiple failure modes. Developing a remaining useful life (RUL) prediction model that does not consider the failure mode causing degradation is likely to result in inaccurate predictions. However, distinguishing between causes of failure without manually inspecting the system is nontrivial. This challenge is increased when the causes of historically observed failures are unknown. Sensors, which are useful for monitoring the state-of-health of systems, can also be used for distinguishing between multiple failure modes as the presence of multiple failure modes results in discriminatory behavior of the sensor signals. When systems are equipped with multiple sensors, some sensors may exhibit behavior correlated with degradation, while other sensors do not. Furthermore, which sensors exhibit this behavior may differ for each failure mode. In this paper, we present a simultaneous clustering and sensor selection approach for unlabeled training datasets of systems exhibiting multiple failure modes. The cluster assignments and the selected sensors are then utilized in real-time to first diagnose the active failure mode and then to predict the system RUL. We validate the methodology using a simulated dataset of systems exhibiting two failure modes and on NASA turbofan degradation dataset.
keywords:
Multiple Failure Modes , Remaining Useful Life , Sensor Fusion , Functional Principal Component Analysis , Mixture of Gaussian Regression , Adaptive Sparse Group Lasso , Expectation Maximization
journal: Reliability Engineering & System Safety
[…]
{highlights}
We propose a two-stage prognostic framework for multi-sensor systems with multiple failure modes.
Offline: An unsupervised method labels failure events and ranks sensor importance.
Online: Sensor features are fused for failure mode diagnosis, then key sensors predict RUL.
Our method achieves $80\%$ clustering accuracy and $<3\%$ RUL prediction error in simulations.
In NASA turbofan data, early predictions outperform benchmarks while reducing input by $30\%$.
1 Introduction
@@
  • Finally, we provide a two-step sequential approach to leverage this knowledge of failure labels and informative sensors to (1) perform active failure mode diagnosis and (2) subsequent RUL prediction on systems degrading in real-time. The paper mentions this as the “online diagnosis and prognostics” step.
The remainder of the paper is outlined as follows. In Section 2, we review the relevant literature associated with this research area. In Section III, we outline our proposed multi-failure mode prognostics framework. In Section 4, we present our offline sensor selection scheme that utilizes an EM algorithm. In Section 5, we present our online diagnostics and prognostics method. It discusses a strategy to fuse informative sensors for real-time diagnosis of failure modes and subsequent prediction of RUL. Our methodology is demonstrated through two case studies – one using a simulated dataset in section 6 and the other NASA turbofan engine dataset in section 7. We conclude by summarizing our framework and key findings in Section 8.
2 Literature Review
@@
3 Overview of the Prognostics Framework
@@
4 Offline Sensor Selection Algorithm
@@
4.1 Feature Extraction: CA-FPCA
@@
4.2 Optimization of the MGR-ASGL model
@@
4.2.1 Mixture of Gaussian Regression (MGR) model
@@
$f_{\vartheta}(y_{i}\mid\boldsymbol{x}_{i,1},\ldots,\boldsymbol{x}_{i,P})=\sum_{k=1}^{K}\pi_{k}\frac{1}{\sqrt{2\pi}\sigma_{k}}\times\ \exp\left(-\frac{1}{2\sigma_{k}^{2}}\left(y_{i}-\beta_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\beta}_{p,k}\right)^{2}\right)$ (3)
4.2.2 Log-Likelihood
@@
$\ell(\boldsymbol{\Theta};\boldsymbol{Y})=\sum_{i=1}^{N}\ln\sum_{k=1}^{K}\pi_{k}\times\ \frac{\rho_{k}}{\sqrt{2\pi}}\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)$ (4)
@@
$\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})=\sum_{i=1}^{N}\sum_{k=1}^{K}\boldsymbol{Z}_{i}[k]\ln\pi_{k}\left(\frac{\rho_{k}}{\sqrt{2\pi}}\right)\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)$ (5)
@@
/* Feature Extraction from all $P$ sensors (CA-FPCA) */
1 for $p=1$ to $P$ do
2       $\mathcal{I}_{k}^{p}\leftarrow$ K-means clustering on FPC-scores;
3       for $k=1$ to $K$ do
4             $\xi_{i,m,p}(k)\leftarrow$ Use $\mathcal{I}_{k}^{p}$ to obtain CA-FPC-scores (Eq. 2);
5
6       end for
7      $x_{i,p}\leftarrow$ Obtain standardized CA-FPC-scores;
8       $q_{p}\leftarrow$ Number of CA-FPC-scores retained for sensor $p$;
9
10 end for
/* MGR-ASGL based EM algorithm */
[…]
12 while $iter\leq iter_{max}$ do
       $\gamma_{i}^{iter}\leftarrow$ Use $x_{i,p}$ in Eq. (7) for all $i$; ;
@@ 8       $q_{p}\leftarrow$ Number of CA-FPC-scores retained for sensor $p$;
       $\Pi^{iter}\leftarrow$ Update using Eq. (9); ;
@@
13       $\boldsymbol{\rho}^{iter},\Psi^{iter}\leftarrow$ Solve Eq. (8) using $\sum_{p=1}^{P}\sqrt{q_{p}}$ and $\gamma_{i}^{iter}$;
14       $iter\leftarrow iter+1$;
15
16 end while
[…]
18 $\boldsymbol{Z}_{i}[\operatorname*{arg\,max}_{k}{\gamma_{i}^{iter_{max}}}]\leftarrow 1$ for all $i$;
19 $\boldsymbol{Z}_{i}[\sim\operatorname*{arg\,max}_{k}{\gamma_{i}^{iter_{max}}}]\leftarrow 0$ for all $i$;
20 foreach $p$ and $k$ do
21       if $\lVert\psi^{iter_{max}}_{p,k}\rVert_{2}\geq\delta$ then
@@ 18 $\boldsymbol{Z}_{i}[\operatorname*{arg\,max}_{k}{\gamma_{i}^{iter_{max}}}]\leftarrow 1$ for all $i$;
4.2.3 Adaptive Sparse Group Lasso (ASGL) penalty
@@
$\min_{\boldsymbol{\Theta}}-\frac{1}{N}\mathbb{E}_{\boldsymbol{\Theta}_{\text{old}}}[\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})|\boldsymbol{Y}]+\lambda\sum_{k=1}^{K}\pi_{k}\sum_{p=1}^{P}\sqrt{q_{p}}\times\left[(1-\alpha)\left\|\boldsymbol{\psi}_{p,k}\right\|_{2}+\alpha\left\|\boldsymbol{\psi}_{p,k}\right\|_{1}\right]$ (6)
@@
4.2.4 Expectation Maximization algorithm
@@
$\min_{\boldsymbol{\Theta}}\frac{1}{N}Q(\boldsymbol{\Theta}\mid\boldsymbol{\Theta}_{\text{old}})+\lambda\sum_{k=1}^{K}\pi_{k}\sum_{p=1}^{P}\sqrt{q_{p}}\times\left[(1-\alpha)\left\|\boldsymbol{\psi}_{p,k}\right\|_{2}+\alpha\left\|\boldsymbol{\psi}_{p,k}\right\|_{1}\right]$ (8)
@@
5 Online diagnosis and Prognostics
@@
3 $\zeta_{(i),h}\leftarrow$ MFPCA on training units (Eq. 10);
4 Train a KNN classifier using $\zeta_{(i),h}$;
5
@@ 4 Train a KNN classifier using $\zeta_{(i),h}$;
6 $\zeta_{test,h}\leftarrow$ MFPCA on test unit;
7 $k^{*}\leftarrow$ Mode of labels in KNN of $\zeta_{test,h}$;
8
@@ 7 $k^{*}\leftarrow$ Mode of labels in KNN of $\zeta_{test,h}$;
9 $\zeta^{k^{*}}_{(i),h}\leftarrow$ MFPCA: selected sensors (Eq. 14);
10 $c_{0}^{k^{*}},\dots,c_{h}^{k^{*}}\leftarrow$ Obtain $\operatorname*{arg\,min}$ in Eq. 17;
11
@@
12 $\zeta^{k^{*}}_{test,h}\leftarrow$ MFPCA on test unit (Eq. 14);
13 $RUL_{test}\leftarrow$ Equation 18;
@@
5.1 Sensor Fusion: MFPCA
@@
5.2 Failure Mode Diagnosis: K-nearest Neighbors
@@
5.3 RUL Prediction: Weighted Functional Regression
@@
5.3.1 Time-Varying Functional Regression
@@
$\min_{c_{0}^{k^{*}},c_{1}^{k^{*}},\ldots,c_{H_{k^{*}}}^{k^{*}}}\sum_{i\in\mathcal{I}_{t^{*}}^{k^{*}}}\left(\sqrt{w_{i}}y_{i}-\sqrt{w_{i}}c_{0}^{k^{*}}-\sum_{h=1}^{H_{k^{*}}}\sqrt{w_{i}}\zeta_{i,h}^{k^{*}}c_{h}^{k^{*}}\right)^{2}+\lambda\sum_{h=1}^{H_{k^{*}}}\lVert c_{h}^{k^{*}}\rVert$ (17)
5.3.2 RUL prediction
@@
6 Case Study 1: Simulated Dataset
@@ 6.1 Simulation Model
Figure 2: Case study 1 data: Signals in sensors 1, 3, 5, 19 (columns 1, 2, 3, 4 respectively) for FM 1 and 2 under different SNRs – [2, 5], [5, 8], and [8, 11] (rows 1, 2, 3 respectively)
6.2 Failure Mode Labeling and Sensor Selection
We analyze the ability of the MGR-ASGL model to accurately cluster the degradation signals into one of the two possible failure modes along with selecting the subset of informative sensors. We start by utilizing CA-FPCA as mentioned in 4.1 to extract features for each sensor separately. To determine the number of CA-FPC scores to retain, we use the 95% fraction of variance explained (FVE) criterion. For each cluster, we select the first $q_{p}^{k}$ CA-FPC-scores whose eigenvalues sum to at least 0.95. Since the number of CA FPC scores to retain can vary for each cluster, we retain the maximum of $q_{p}^{k}$ over both failure modes. For the remainder of this paper, we use 95% FVE criterion for determining the number of FPC scores, CA-FPC scores, and MFPC scores to retain. The CA-FPC scores serve as inputs to the MGR-ASGL model. Since this model has two tuning parameters, we utilize 5-fold cross-validation with $\lambda\in\{0.0050,0.0258,0.0466,...,0.3792,0.4000\}$ and $\alpha\in\{0.00,0.25,0.50,0.75,1.00\}$. We use minimum mean-squared error (minMSE) criterion, which selects the tuning parameters that minimize the mean-squared prediction error averaged over all folds. Table 1 displays the tuning parameter selection using both criteria over all three datasets. For each dataset and criterion, we fit the MGR-ASGL model using these tuning parameters.
Table 1: Tuning parameter selection for case study 1
@@ 6.2 Failure Mode Labeling and Sensor Selection
We then evaluate the performance of the MGR-ASGL model on the three datasets. While clustering is an unsupervised learning task, we know the ground truth regarding the failure modes. Therefore, we evaluate the ability of the clustering algorithm to match the ground truth. Furthermore, we have predetermined which sensors are most informative and we would like to determine whether or not the MGR-ASGL selects these sensors. The results of this performance evaluation are shown in Table 2.
Table 2: MGR-ASGL performance for case study 1
@@
For all scenarios, the initial failure mode labels were determined randomly. Table II shows that for the lowest and the highest signal-to-noise ratios, the clustering algorithm was capable of achieving at worst 78.8% accuracy for a particular failure mode. The exception is the dataset [5,8], where the algorithm has low accuracy at clustering. This is due to the tuning parameter combination resulting in minimal coefficient shrinkage that enabled the nonsignificant sensors to influence cluster assignments. A slightly larger $\lambda$ value would decrease this error. For all three models, several sensors that are deemed ”noninformative” were selected. This can be attributed to some of these sensors having high values of $\rho_{p}^{k}$ relative to the specified range of [0.1,0.6]. However, simply reporting the sensors selected does not account for how informative they are. Table 3 reports the four sensors with the highest $l_{2}$-norm of the regression coefficients associated with that sensor. Except for the [5,8] dataset, at least three of the four sensors with the largest $l_{2}$-norm for each failure mode are elements of the set of informative sensors. This indicates that the MGR-ASGL model results in the informative sensors having larger regression coefficients (in magnitude) than the noninformative sensors. Given the selected sensors, we want to test the robustness of the online portion of the methodology to predict remaining life.
Table 3: $l_{2}$-norms of four most informative sensors for FM1 & FM2